# The star HIP 41378 potentially misaligned with its cohort of long-period planets

### v1

@@ 4.1 A likely misaligned planetary system
The presence of a massive, inclined outer planet can tilt inner planets with nodal precession operating on a million-year timescale. This mechanism has been proposed to explain the misalignment of HAT-P-11 b, for example (Yee et al. 2018). However, with seven years of RV monitoring (Santerne et al. 2019, Grouffal et al., in prep.), we detected no signs of a very long-period planet more massive than 0.15 at a period longer than 2000 days. We, therefore, excluded perturbation from another external planetary companion to explain the tilt of HIP 41378 d and f.
@@ 4.2 Resonances and spin-orbit angles
Beyond the measured spin-orbit misalignment, any theory aiming to explain the HIP 41378 system’s architecture must also account for its multiplicity and resonant structure. The two inner planets, HIP 41378 b and c, orbit at 15.6 and 31.7 days, respectively, placing them near a 2:1 mean motion resonance (MMR) (Vanderburg et al. 2016). The outer planets lie in or close to a 4:3–3:2 resonant chain (Santerne et al. 2019), suggesting an ordered orbital configuration. A system in resonant configuration has also been proposed to explain the unusually low density of HIP 41378 f, potentially linked to the presence of circumplanetary rings (Lu et al. 2025). Resonant configurations are generally associated with smooth formation processes, making it unlikely that violent dynamical interactions contributed to the observed misalignment (Rice et al. 2023). This strengthens the case for a primordial origin of the system’s tilt, as dynamical processes capable of producing large obliquities typically disrupt resonances (Izidoro et al. 2017; Esteves et al. 2023). HIP 41378 is not the only multi-planet system known to exhibit misalignment. In the Kepler-56 system, two close-in, coplanar planets with orbital periods of 10.50 and 21.41 days were found to be misaligned by approximately 45∘, based on a combination of asteroseismology and dynamical modelling (Huber et al. 2013). The detection of a distant outer companion in that system supports the hypothesis that perturbations from an inclined outer planet may have driven the inner planets’ tilt. A similar scenario has been proposed for Kepler-129, a system composed of two planets at 15.79 and 82.20 days. There, a stellar inclination of 38∘ has been inferred, suggesting a misalignment that may also result from the influence of a massive companion. Another case is HD 3167, a system in which the innermost planet is nearly aligned with the stellar rotation axis, while the two outer planets (c and d) are found to orbit in a plane almost perpendicular to that of the inner planet (Dalal et al. 2019; Bourrier et al. 2021). As discussed in Grouffal et al. (2022), this configuration has a very low probability for the HIP 41378’s planetary system considering the low mutual inclination between the planets and the long periods of the planets. Figure 6 summarises the current sample of confirmed exoplanets with measured obliquities. The low projected obliquity of HIP 41378 f is consistent with the findings of Albrecht et al. (2013), which suggest that planets in multi-planet systems tend to be well-aligned. However, the measured 3D obliquity of HIP 41378 f deviates from previously observed values in such systems. To date, only a handful of multi-planet systems have obliquity measurements, and these are only for planets with orbital periods shorter than 50 days. Additional observations of long-period planets, particularly using a combination of instruments, would help to better interpret the measurement for HIP 41378 f and provide further constraints on the formation and evolutionary pathways of long-period exoplanets. In particular, comparing more systems with planets near or far from mean-motion resonances is needed to determine whether different evolutionary pathways are at play.
@@ 5 Conclusions
The RM effect of HIP 41378 f was significantly detected during its transit on 12–13 November 2022. Our analysis suggests that the planet orbits the star in a prograde direction, potentially misaligned with the stellar spin axis. Our results show that HIP 41378 f has a slightly misaligned projected obliquity of ∘. Although all analyses converge toward a projected obliquity between 10 and 30 degrees, the egress part of the transit is missing. It is therefore possible that the derived uncertainties are overly optimistic. However, even in the case of a projected obliquity close to zero, the values of obtained from two independent analyses using starry, including differential rotation, and ARoME, including macroturbulence, are inconsistent with a stellar inclination of . Both the photometric and RV analyses converge toward a stellar rotation period between 6 and 9 days. A longer rotation period of 11 days, which would imply a nearly aligned system, is not supported by any dataset. The three-dimensional obliquity of the system is non-zero. However, we acknowledge that this conclusion relies on the measured stellar rotation period. Combining the observed RM signal with precise stellar parameters, we determined a 3D spin-orbit angle of ∘.
@@
Looking ahead, the upcoming PLATO mission (Rauer et al. 2014), which aims to uncover and characterise long-period planets, will offer a unique opportunity to validate and refine theories of planetary formation in diverse environments. International collaboration like for HIP 41378 f will help to follow-up systems that PLATO will discover.
@@ Appendix A Stellar rotation period analysis
Appendix B Summary of priors and posteriors for the RM-transit Analysis
@@
Appendix C ARoME analysis with macroturbulences

### v2

@@ 4.1 A likely misaligned planetary system
The presence of a massive, inclined outer planet can tilt inner planets with nodal precession operating on a million-year timescale. This mechanism has been proposed to explain the misalignment of HAT-P-11 b, for example (Yee et al. 2018). However, with seven years of RV monitoring (Santerne et al. 2019, Grouffal et al. in prep.), we detected no signs of a very long-period planet more massive than 0.15 at a period longer than 2000 days. We, therefore, excluded perturbation from another external planetary companion to explain the tilt of HIP 41378 d and f.
@@ 4.2 Resonances and spin-orbit angles
Beyond the measured spin-orbit misalignment, any theory aiming to explain the HIP 41378 system’s architecture must also account for its multiplicity and resonant structure. The two inner planets, HIP 41378 b and c, orbit at 15.6 and 31.7 days, respectively, placing them near a 2:1 mean motion resonance (MMR) (Vanderburg et al. 2016). The outer planets lie in or close to a 4:3-3:2 resonant chain (Santerne et al. 2019), suggesting an ordered orbital configuration. A system in resonant configuration has also been proposed to explain the unusually low density of HIP 41378 f, potentially linked to the presence of circumplanetary rings (Lu et al. 2025). Resonant configurations are generally associated with smooth formation processes, making it unlikely that violent dynamical interactions contributed to the observed misalignment (Rice et al. 2023). This strengthens the case for a primordial origin of the system’s tilt, as dynamical processes capable of producing large obliquities typically disrupt resonances (Izidoro et al. 2017; Esteves et al. 2023). HIP 41378 is not the only multi-planet system known to exhibit misalignment. In the Kepler-56 system, two close-in, coplanar planets with orbital periods of 10.50 and 21.41 days were found to be misaligned by approximately 45∘, based on a combination of asteroseismology and dynamical modelling (Huber et al. 2013). The detection of a distant outer companion in that system supports the hypothesis that perturbations from an inclined outer planet may have driven the inner planets’ tilt. A similar scenario has been proposed for Kepler-129, a system composed of two planets at 15.79 and 82.20 days. There, a stellar inclination of 38∘ has been inferred, suggesting a misalignment that may also result from the influence of a massive companion. Another case is HD 3167, a system in which the innermost planet is nearly aligned with the stellar rotation axis, while the two outer planets (c and d) are found to orbit in a plane almost perpendicular to that of the inner planet (Dalal et al. 2019; Bourrier et al. 2021). As discussed in Grouffal et al. (2022), this configuration has a very low probability for the HIP 41378’s planetary system considering the low mutual inclination between the planets and the long periods of the planets. Figure 6 summarises the current sample of confirmed exoplanets with measured obliquities. The low projected obliquity of HIP 41378 f is consistent with the findings of Albrecht et al. (2013), which suggest that planets in multi-planet systems tend to be well-aligned. However, the measured 3D obliquity of HIP 41378 f deviates from previously observed values in such systems. To date, only a handful of multi-planet systems have obliquity measurements, and these are only for planets with orbital periods shorter than 50 days. Additional observations of long-period planets, particularly using a combination of instruments, would help to better interpret the measurement for HIP 41378 f and provide further constraints on the formation and evolutionary pathways of long-period exoplanets. In particular, comparing more systems with planets near or far from mean-motion resonances is needed to determine whether different evolutionary pathways are at play.
@@ 5 Conclusions
The RM effect of HIP 41378 f was significantly detected during its transit on 12–13 November 2022. Our analysis indicates that the planet follows a prograde orbit, potentially misaligned with the stellar spin axis. Our results show that HIP 41378 f has a slightly misaligned projected obliquity of ∘. Although all analyses converge toward a projected obliquity between 10 and 30 degrees, the egress part of the transit is missing. It is therefore possible that the derived uncertainties are overly optimistic. However, even in the case of a projected obliquity close to zero, the values of obtained from two independent analyses using starry, including differential rotation, and ARoME, including macroturbulence, are inconsistent with a stellar inclination of . Both the photometric and RV analyses converge toward a stellar rotation period between 6 and 9 days. A longer rotation period of 11 days, which would imply a nearly aligned system, is not supported by any dataset. The three-dimensional obliquity of the system is non-zero. However, we acknowledge that this conclusion relies on the measured stellar rotation period. Combining the observed RM signal with precise stellar parameters, we determined a 3D spin-orbit angle of ∘.
@@
Looking ahead, the upcoming PLATO mission (Rauer et al. 2014), which aims to uncover and characterise long-period planets, will offer a unique opportunity to validate and refine theories of planetary formation in diverse environments. International collaboration, like for HIP 41378 f, will help to follow-up systems that PLATO will discover.
@@ Appendix A Stellar rotation period analysis
Appendix B ESPRESSO out-of-transit CCF
Appendix C Summary of the priors and posteriors for the RM-transit Analysis
@@
Appendix D ARoME analysis with macroturbulences